# Quadratic Gating Functions in Mixture of Experts: A Statistical Insight

### v2

@@ Appendix D Proof of Theoretical Results
D.1 Proof of Theorem 3.1
For the proof of the theorem, we first introduce some notation. Firstly, we denote by $\mathcal{F}_{N}(\Theta)$ the set of conditional densities of all mixing measures in $\mathcal{G}_{N}(\Theta)$, that is, $\mathcal{F}_{N}(\Theta):=\{f_{G}({\bm{x}}):G\in\mathcal{G}_{N}(\Theta)\}$. Additionally, for each $\delta>0$, the $L_{2}(\mu)$ ball centered around the regression function $f_{G_{*}}({\bm{x}})$ and intersected with the set $\mathcal{F}_{N}(\Theta)$ is defined as
@@
In order to measure the size of the above set, Geer et al. [40] suggest using the following quantity:
@@
where $H_{B}(t,\mathcal{F}_{N}(\Theta,t),\|\cdot\|_{L^{2}(\mu))}$ stands for the bracketing entropy [40] of $\mathcal{F}_{N}(\Theta,u)$ under the $L^{2}$-norm, and $t\vee\delta:=\max\{t,\delta\}$. By using the similar proof argument of Theorem 7.4 and Theorem 9.2 in [40] with notations being adapted to this work, we obtain the following lemma:
Lemma D.1.
@@
For each mixing measure $G=\sum_{i=1}^{N}\exp(c_{i})\delta_{({\bm{A}}_{i},{\bm{b}}_{i},{\bm{\eta}}_{i})}\in\mathcal{G}_{N}(\Theta)$, we consider other two mixing measures:
@@
$\|f_{G}-f_{\check{G}}\|_{L^{\infty}}$ $=\Bigg{\|}\sum_{i=1}^{N}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}\cdot[h({\bm{x}},{\bm{\eta}}_{i})-h({\bm{x}},\overline{{\bm{\eta}}}_{i})]\Bigg{\|}_{L^{\infty}}$
$\leq\sum_{i=1}^{N}\Bigg{\|}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}\cdot[h({\bm{x}},{\bm{\eta}}_{i})-h({\bm{x}},\overline{{\bm{\eta}}}_{i})]\Bigg{\|}_{L^{\infty}}$
$\leq\sum_{i=1}^{N}\|h({\bm{x}},{\bm{\eta}}_{i})-h({\bm{x}},\overline{{\bm{\eta}}}_{i})\|_{L^{\infty}}$
@@
Here, the first inequality is according to the triangle inequality, the second inequality occurs as the softmax weight is bounded by 1, and the third inequality follows from the fact that the expert $h({\bm{x}},\cdot)$ is a Lipschitz function. Next, we have
$\|f_{\check{G}}-f_{\overline{G}}\|_{L^{\infty}}$ $=\Bigg{\|}\sum_{i=1}^{N}\Bigg{[}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}-\frac{\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{i}{\bm{x}}+(\overline{{\bm{b}}}_{i})^{\top}{\bm{x}}+\overline{c}_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{j}{\bm{x}}+(\overline{{\bm{b}}}_{j})^{\top}{\bm{x}}+\overline{c}_{j})}\Bigg{]}\cdot h({\bm{x}},\overline{{\bm{\eta}}}_{i})\Bigg{\|}_{L^{\infty}}$
$\leq\sum_{i=1}^{N}\Bigg{\|}\Bigg{[}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}-\frac{\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{i}{\bm{x}}+(\overline{{\bm{b}}}_{i})^{\top}{\bm{x}}+\overline{c}_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{j}{\bm{x}}+(\overline{{\bm{b}}}_{j})^{\top}{\bm{x}}+\overline{c}_{j})}\Bigg{]}\cdot h({\bm{x}},\overline{{\bm{\eta}}}_{i})\Bigg{\|}_{L^{\infty}}$
$\leq\sum_{i=1}^{N}\Bigg{\|}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}-\frac{\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{i}{\bm{x}}+(\overline{{\bm{b}}}_{i})^{\top}{\bm{x}}+\overline{c}_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{j}{\bm{x}}+(\overline{{\bm{b}}}_{j})^{\top}{\bm{x}}+\overline{c}_{j})}\Bigg{\|}_{L^{\infty}}$
$\lesssim\sum_{i=1}^{N}\Big{[}\|{\bm{A}}_{i}-\overline{{\bm{A}}}_{i}\|\cdot\|{\bm{x}}\|^{2}+\|{\bm{b}}_{i}-\overline{{\bm{b}}}_{i}\|\cdot\|{\bm{x}}\|+|c_{i}-\overline{c}_{i}|\Big{]}$
@@
Let $\Psi(\delta)=\delta\cdot[\log(1/\delta)]^{1/2}$, then $\Psi(\delta)/\delta^{2}$ is a non-increasing function of $\delta$. Furthermore, equation (32) indicates that $\Psi(\delta)\geq\mathcal{J}_{B}(\delta,\mathcal{F}_{N}(\Theta,\delta))$. In addition, let $\delta_{n}=\sqrt{\log(n)/n}$, then we get that $\sqrt{n}\delta^{2}_{n}\geq c\Psi(\delta_{n})$ for some universal constant $c$. Finally, by applying Lemma D.1, we achieve the desired conclusion of the theorem.
D.2 Proof of Theorem 3.3
@@
$\inf_{G\in\mathcal{G}_{N}(\Theta)}\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{L}_{1}(G,G_{*})>0.$ (33)
@@
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{G}_{N}(\Theta):\mathcal{L}_{1}(G,G_{*})\leq\varepsilon}\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{L}_{1}(G,G_{*})>0.$ (34)
Assume by contrary that the above inequality does not hold true, then there exists a sequence of mixing measures $G_{n}=\sum_{i=1}^{N^{*}}\exp(c_{i}^{n})\delta_{({\bm{A}}^{n}_{i},{\bm{b}}^{n}_{i},{\bm{\eta}}^{n}_{i})}$ in $\mathcal{G}_{N}(\Theta)$ such that $\mathcal{L}_{1n}:=\mathcal{L}_{1}(G_{n},G_{*})\to 0$ and
@@
Decomposition of $A_{n}$. Next, we continue to separate the term $A_{n}$ into two parts as follows:
$A_{n}({\bm{x}})$ $:=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})\Big{[}\exp({\bm{x}}^{\top}A_{i}^{n}{\bm{x}}+(b_{i}^{n})^{\top}{\bm{x}})h({\bm{x}};\eta_{i}^{n})-\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}}+({\bm{b}}_{j}^{*})^{\top}{\bm{x}})h({\bm{x}};{\bm{\eta}}_{j}^{*})\Big{]}$
$+\sum_{j:|\mathcal{V}_{j}|>1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})\Big{[}\exp({\bm{x}}^{\top}A_{i}^{n}{\bm{x}}+(b_{i}^{n})^{\top}{\bm{x}})h({\bm{x}};\eta_{i}^{n})-\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}}+({\bm{b}}_{j}^{*})^{\top}{\bm{x}})h({\bm{x}};{\bm{\eta}}_{j}^{*})\Big{]}$
@@
×$\hskip 142.26378pt\times\frac{\partial^{|\alpha_{1}|+|\alpha_{2}|}E}{\partial A^{\alpha_{1}}\partial b^{\alpha_{2}}}({\bm{x}};{\bm{A}}_{j}^{*},{\bm{b}}_{j}^{*})\frac{\partial^{|\alpha_{3}|}h}{\partial{\bm{\eta}}^{\alpha_{3}}}({\bm{x}};{\bm{\eta}}_{j}^{*})+R_{n,1}({\bm{x}}),$
@@
×$\hskip 142.26378pt\times\frac{\partial^{2|\alpha_{1}|+|\alpha_{2}|}E}{\partial b^{\tau(\alpha_{1},\alpha_{2})}}({\bm{x}};{\bm{A}}_{j}^{*},{\bm{b}}_{j}^{*})\frac{\partial^{|\alpha_{3}|}h}{\partial{\bm{\eta}}^{\alpha_{3}}}({\bm{x}};{\bm{\eta}}_{j}^{*})+R_{n,1}({\bm{x}})$
$=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{|\alpha_{3}|=0}^{1}\sum_{|\ell_{1}|=0\vee 1-|\alpha_{3}|}^{2(1-|\alpha_{3}|)}S_{n,j,\alpha_{3},\ell_{1}}\cdot\frac{\partial^{|\ell_{1}|}E}{\partial b^{\ell_{1}}}({\bm{x}};{\bm{A}}_{j}^{*},{\bm{b}}_{j}^{*})\frac{\partial^{|\alpha_{3}|}h}{\partial{\bm{\eta}}^{\alpha_{3}}}({\bm{x}};{\bm{\eta}}_{j}^{*})+R_{n,1}({\bm{x}}),$
@@
$A_{n,2}({\bm{x}})$ $=\sum_{j:|\mathcal{V}_{j}|>1}\sum_{|\alpha_{3}|=0}^{\bar{r}_{j}}\sum_{|\ell_{1}|=0\vee 1-|\alpha_{3}|}^{2(\bar{r}_{j}-|\alpha_{3}|)}S_{n,j,\alpha_{3},\ell_{1}}\cdot\frac{\partial^{|\ell_{1}|}E}{\partial b^{\ell_{1}}}({\bm{x}};{\bm{A}}_{j}^{*},{\bm{b}}_{j}^{*})\frac{\partial^{|\alpha_{3}|}h}{\partial{\bm{\eta}}^{\alpha_{3}}}({\bm{x}};{\bm{\eta}}_{j}^{*})+R_{n,2}({\bm{x}}),$
@@
Decomposition of $B_{n}$. Note that $B_{n}({\bm{x}})$ can be rewritten as
@@
$Q_{n}({\bm{x}})$ $=\sum_{j=1}^{N^{*}}\sum_{|\alpha_{3}|=0}^{\bar{r}_{j}}\sum_{|\ell_{1}|=0}^{2(\bar{r}_{j}-|\alpha_{3}|)}S_{n,j,\alpha_{3},\ell_{1}}\cdot\frac{\partial^{|\ell_{1}|}E}{\partial b^{\ell_{1}}}({\bm{x}};{\bm{A}}_{j}^{*},{\bm{b}}_{j}^{*})\frac{\partial^{|\alpha_{3}|}h}{\partial{\bm{\eta}}^{\alpha_{3}}}({\bm{x}};{\bm{\eta}}_{j}^{*})$
@@
as $n\to\infty$. It follows from the limits in equations (40), (44), (45) and the formulation of $\mathcal{L}_{1n}$ in equation (D.2) that
@@
Subsequently, we define ${M}_{n}:=\max\{|(\Delta A_{i1}^{n})^{(11)}|^{1/2},|(\Delta b_{i1}^{n})^{(1)}|:i\in\mathcal{V}_{1}\}$ and ${\pi}_{n}=\max_{i\in\mathcal{V}_{1}}\exp(c_{i}^{n})$. For any $i\in\mathcal{V}_{1}$, it is clear that the sequence of positive real numbers $(\exp(c_{i}^{n})/{\pi}_{n})$ is bounded, therefore, we can replace it by its subsequence that admits a non-negative limit denoted by $p_{i}^{2}=\lim_{n\to\infty}\exp(c_{i}^{n})/{\pi}_{n}$. In addition, let us denote $(\Delta A_{i1}^{n})^{(11)}/{M}_{n}^{2}\to\gamma_{1i}$ and $(\Delta b_{i1}^{n})^{(1)}/{M}_{n}\to\gamma_{2i}$. Since $\exp(c_{i}^{n})\geq\beta$ for some $\beta>0$, the real numbers $p_{i}$ will not vanish, and at least one of them is equal to $1$. Analogously, at least one of the terms $\gamma_{1i}$ and $\gamma_{2i}$ is equal to either $1$ or $-1$.
@@
with a note that at least one among them is non-zero. Then, from the decomposition of $Q_{n}({\bm{x}})$ in equation (D.2), we have
$\sum_{j=1}^{N^{*}}\sum_{|\alpha_{3}|=0}^{\bar{r}_{j}}\sum_{|\ell_{1}|=0}^{2(\bar{r}_{j}-|\alpha_{3}|)}\phi_{j,\alpha_{3},\ell_{1}}\cdot$ $\frac{\partial^{|\ell_{1}|}E}{\partial b^{\ell_{1}}}({\bm{x}};{\bm{A}}_{j}^{*},{\bm{b}}_{j}^{*})\frac{\partial^{|\alpha_{3}|}h}{\partial{\bm{\eta}}^{\alpha_{3}}}({\bm{x}};{\bm{\eta}}_{j}^{*})$
@@
$\inf_{G\in\mathcal{G}_{N}(\Theta):\mathcal{L}_{1}(G,G_{*})\leq\varepsilon^{\prime}}\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{L}_{1}(G,G_{*})>0.$
@@
$\inf_{G\in\mathcal{G}_{N}(\Theta):\mathcal{L}_{1}(G,G_{*})>\varepsilon^{\prime}}\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{L}_{1}(G,G_{*})>0.$ (53)
Assume by contrary that the inequality (53) does not hold true, then we can find a sequence of mixing measures $G^{\prime}_{n}\in\mathcal{G}_{N}(\Theta)$ such that $\mathcal{L}_{1}(G^{\prime}_{n},G_{*})>\varepsilon^{\prime}$ and

### v3

@@ Appendix A Proof of Theoretical Results
A.1 Proof of Theorem 3.1
For the proof of the theorem, we first introduce some notation. Firstly, we denote by $\mathcal{F}_{N}(\Theta)$ the set of conditional densities of all mixing measures in $\mathcal{M}_{N}(\Theta)$, that is, $\mathcal{F}_{N}(\Theta):=\{f_{G}({\bm{x}}):G\in\mathcal{M}_{N}(\Theta)\}$. Additionally, for each $\delta>0$, the $L_{2}(\mu)$ ball centered around the regression function $f_{G_{*}}({\bm{x}})$ and intersected with the set $\mathcal{F}_{N}(\Theta)$ is defined as
@@
In order to measure the size of the above set, Geer et al. vandeGeer-00 suggest using the following quantity:
@@
where $H_{B}(t,\mathcal{F}_{N}(\Theta,t),\|\cdot\|_{L^{2}(\mu))}$ stands for the bracketing entropy vandeGeer-00 of $\mathcal{F}_{N}(\Theta,u)$ under the $L^{2}$-norm, and $t\vee\delta:=\max\{t,\delta\}$. By using the similar proof argument of Theorem 7.4 and Theorem 9.2 in vandeGeer-00 with notations being adapted to this work, we obtain the following lemma:
Lemma A.1.
@@
For each mixing measure $G=\sum_{i=1}^{N}\exp(c_{i})\delta_{({\bm{A}}_{i},{\bm{b}}_{i},{\bm{\eta}}_{i})}\in\mathcal{M}_{N}(\Theta)$, we consider other two mixing measures:
@@
$\|f_{G}-f_{\check{G}}\|_{L^{\infty}}$ $=\sup_{{\bm{x}}\in\mathcal{X}}\Bigg{|}\sum_{i=1}^{N}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{N}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}\cdot[\mathcal{E}({\bm{x}},{\bm{\eta}}_{i})-\mathcal{E}({\bm{x}},\overline{{\bm{\eta}}}_{i})]\Bigg{|}$
$\leq\sum_{i=1}^{N}\sup_{{\bm{x}}\in\mathcal{X}}\Bigg{|}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{N}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}\cdot[\mathcal{E}({\bm{x}},{\bm{\eta}}_{i})-\mathcal{E}({\bm{x}},\overline{{\bm{\eta}}}_{i})]\Bigg{|}$
$\leq\sum_{i=1}^{N}\sup_{{\bm{x}}\in\mathcal{X}}|\mathcal{E}({\bm{x}},{\bm{\eta}}_{i})-\mathcal{E}({\bm{x}},\overline{{\bm{\eta}}}_{i})|$
@@
Here, the first inequality is according to the triangle inequality, the second inequality occurs as the softmax weight is bounded by 1, and the third inequality follows from the fact that the expert $\mathcal{E}({\bm{x}},\cdot)$ is a Lipschitz function. Next, we have
$\|f_{\check{G}}-f_{\overline{G}}\|_{L^{\infty}}$
$=\sup_{{\bm{x}}\in\mathcal{X}}\Bigg{|}\sum_{i=1}^{N}\Bigg{[}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{N}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}-\frac{\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{i}{\bm{x}}+(\overline{{\bm{b}}}_{i})^{\top}{\bm{x}}+\overline{c}_{i})}{\sum_{j=1}^{N}\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{j}{\bm{x}}+(\overline{{\bm{b}}}_{j})^{\top}{\bm{x}}+\overline{c}_{j})}\Bigg{]}\mathcal{E}({\bm{x}},\overline{{\bm{\eta}}}_{i})\Bigg{|}$
$\leq\sum_{i=1}^{N}\sup_{{\bm{x}}\in\mathcal{X}}\Bigg{|}\Bigg{[}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{N}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}-\frac{\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{i}{\bm{x}}+(\overline{{\bm{b}}}_{i})^{\top}{\bm{x}}+\overline{c}_{i})}{\sum_{j=1}^{N}\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{j}{\bm{x}}+(\overline{{\bm{b}}}_{j})^{\top}{\bm{x}}+\overline{c}_{j})}\Bigg{]}\mathcal{E}({\bm{x}},\overline{{\bm{\eta}}}_{i})\Bigg{|}$
$\leq\sum_{i=1}^{N}\sup_{{\bm{x}}\in\mathcal{X}}\Bigg{|}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{N}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}-\frac{\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{i}{\bm{x}}+(\overline{{\bm{b}}}_{i})^{\top}{\bm{x}}+\overline{c}_{i})}{\sum_{j=1}^{N}\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{j}{\bm{x}}+(\overline{{\bm{b}}}_{j})^{\top}{\bm{x}}+\overline{c}_{j})}\Bigg{|}$
$\lesssim\sum_{i=1}^{N}\sup_{{\bm{x}}\in\mathcal{X}}\Big{[}\|{\bm{A}}_{i}-\overline{{\bm{A}}}_{i}\|\cdot\|{\bm{x}}\|^{2}+\|{\bm{b}}_{i}-\overline{{\bm{b}}}_{i}\|\cdot\|{\bm{x}}\|+|c_{i}-\overline{c}_{i}|\Big{]}$
@@
Let $\Psi(\delta)=\delta\cdot[\log(1/\delta)]^{1/2}$, then $\Psi(\delta)/\delta^{2}$ is a non-increasing function of $\delta$. Furthermore, equation (31) indicates that $\Psi(\delta)\geq\mathcal{J}_{B}(\delta,\mathcal{F}_{N}(\Theta,\delta))$. In addition, let $\delta_{n}=\sqrt{\log(n)/n}$, then we get that $\sqrt{n}\delta^{2}_{n}\geq c\Psi(\delta_{n})$ for some universal constant $c$. Finally, by applying Lemma A.1, we achieve the desired conclusion of the theorem.
A.2 Proof of Theorem 3.3
@@
$\inf_{G\in\mathcal{M}_{N}(\Theta)}\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{L}_{1}(G,G_{*})>0.$ (32)
@@
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{M}_{N}(\Theta):\mathcal{L}_{1}(G,G_{*})\leq\varepsilon}\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{L}_{1}(G,G_{*})>0.$ (33)
Assume by contrary that the above inequality does not hold true, then there exists a sequence of mixing measures $G_{n}=\sum_{i=1}^{N^{*}}\exp(c_{i}^{n})\delta_{({\bm{A}}^{n}_{i},{\bm{b}}^{n}_{i},{\bm{\eta}}^{n}_{i})}$ in $\mathcal{M}_{N}(\Theta)$ such that $\mathcal{L}_{1n}:=\mathcal{L}_{1}(G_{n},G_{*})\to 0$ and
@@
Decomposition of $A_{n}({\bm{x}})$. Next, we continue to separate the term $A_{n}({\bm{x}})$ into two parts as follows:
$A_{n}({\bm{x}})$
$:=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})\Big{[}\exp({\bm{x}}^{\top}A_{i}^{n}{\bm{x}}+(b_{i}^{n})^{\top}{\bm{x}})\mathcal{E}({\bm{x}};\eta_{i}^{n})-\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}}+({\bm{b}}_{j}^{*})^{\top}{\bm{x}})\mathcal{E}({\bm{x}};{\bm{\eta}}_{j}^{*})\Big{]}$
$+\sum_{j:|\mathcal{V}_{j}|>1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})\Big{[}\exp({\bm{x}}^{\top}A_{i}^{n}{\bm{x}}+(b_{i}^{n})^{\top}{\bm{x}})\mathcal{E}({\bm{x}};\eta_{i}^{n})-\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}}+({\bm{b}}_{j}^{*})^{\top}{\bm{x}})\mathcal{E}({\bm{x}};{\bm{\eta}}_{j}^{*})\Big{]}$
@@
×$\hskip 142.26378pt\times\frac{\partial^{|\alpha_{1}|+|\alpha_{2}|}E}{\partial A^{\alpha_{1}}\partial b^{\alpha_{2}}}({\bm{x}};{\bm{A}}_{j}^{*},{\bm{b}}_{j}^{*})\frac{\partial^{|\alpha_{3}|}\mathcal{E}}{\partial{\bm{\eta}}^{\alpha_{3}}}({\bm{x}};{\bm{\eta}}_{j}^{*})+R_{n,1}({\bm{x}}),$
@@
×$\hskip 142.26378pt\times\frac{\partial^{2|\alpha_{1}|+|\alpha_{2}|}E}{\partial b^{\tau(\alpha_{1},\alpha_{2})}}({\bm{x}};{\bm{A}}_{j}^{*},{\bm{b}}_{j}^{*})\frac{\partial^{|\alpha_{3}|}\mathcal{E}}{\partial{\bm{\eta}}^{\alpha_{3}}}({\bm{x}};{\bm{\eta}}_{j}^{*})+R_{n,1}({\bm{x}})$
$=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{|\alpha_{3}|=0}^{1}\sum_{|\ell_{1}|=0\vee 1-|\alpha_{3}|}^{2(1-|\alpha_{3}|)}S_{n,j,\alpha_{3},\ell_{1}}\cdot\frac{\partial^{|\ell_{1}|}E}{\partial b^{\ell_{1}}}({\bm{x}};{\bm{A}}_{j}^{*},{\bm{b}}_{j}^{*})\frac{\partial^{|\alpha_{3}|}\mathcal{E}}{\partial{\bm{\eta}}^{\alpha_{3}}}({\bm{x}};{\bm{\eta}}_{j}^{*})+R_{n,1}({\bm{x}}),$
@@
$A_{n,2}({\bm{x}})$ $=\sum_{j:|\mathcal{V}_{j}|>1}\sum_{|\alpha_{3}|=0}^{\bar{r}_{j}}\sum_{|\ell_{1}|=0\vee 1-|\alpha_{3}|}^{2(\bar{r}_{j}-|\alpha_{3}|)}S_{n,j,\alpha_{3},\ell_{1}}\cdot\frac{\partial^{|\ell_{1}|}E}{\partial b^{\ell_{1}}}({\bm{x}};{\bm{A}}_{j}^{*},{\bm{b}}_{j}^{*})\frac{\partial^{|\alpha_{3}|}\mathcal{E}}{\partial{\bm{\eta}}^{\alpha_{3}}}({\bm{x}};{\bm{\eta}}_{j}^{*})+R_{n,2}({\bm{x}}),$
@@
Decomposition of $B_{n}({\bm{x}})$. Note that $B_{n}({\bm{x}})$ can be rewritten as
@@
$Q_{n}({\bm{x}})$ $=\sum_{j=1}^{N^{*}}\sum_{|\alpha_{3}|=0}^{\bar{r}_{j}}\sum_{|\ell_{1}|=0}^{2(\bar{r}_{j}-|\alpha_{3}|)}S_{n,j,\alpha_{3},\ell_{1}}\cdot\frac{\partial^{|\ell_{1}|}E}{\partial b^{\ell_{1}}}({\bm{x}};{\bm{A}}_{j}^{*},{\bm{b}}_{j}^{*})\frac{\partial^{|\alpha_{3}|}\mathcal{E}}{\partial{\bm{\eta}}^{\alpha_{3}}}({\bm{x}};{\bm{\eta}}_{j}^{*})$
@@
as $n\to\infty$. It follows from the limits in equations (39), (43), (44) and the formulation of $\mathcal{L}_{1n}$ in equation (A.2) that
@@
Subsequently, we define
${M}_{n}$ $:=\max\{|(\Delta A_{i1}^{n})^{(11)}|^{1/2},|(\Delta b_{i1}^{n})^{(1)}|:i\in\mathcal{V}_{1}\},$
${\pi}_{n}$ $:=\max_{i\in\mathcal{V}_{1}}\exp(c_{i}^{n}).$
For any $i\in\mathcal{V}_{1}$, it is clear that the sequence of positive real numbers $(\exp(c_{i}^{n})/{\pi}_{n})$ is bounded, therefore, we can replace it by its subsequence that admits a non-negative limit denoted by $p_{i}^{2}=\lim_{n\to\infty}\exp(c_{i}^{n})/{\pi}_{n}$. In addition, let us denote $(\Delta A_{i1}^{n})^{(11)}/{M}_{n}^{2}\to\gamma_{1i}$ and $(\Delta b_{i1}^{n})^{(1)}/{M}_{n}\to\gamma_{2i}$. Since $\exp(c_{i}^{n})\geq\beta$ for some $\beta>0$, the real numbers $p_{i}$ will not vanish, and at least one of them is equal to $1$. Analogously, at least one of the terms $\gamma_{1i}$ and $\gamma_{2i}$ is equal to either $1$ or $-1$.
@@
with a note that at least one among them is non-zero. Then, from the decomposition of $Q_{n}({\bm{x}})$ in equation (A.2), we have
$\sum_{j=1}^{N^{*}}\sum_{|\alpha_{3}|=0}^{\bar{r}_{j}}\sum_{|\ell_{1}|=0}^{2(\bar{r}_{j}-|\alpha_{3}|)}\phi_{j,\alpha_{3},\ell_{1}}\cdot$ $\frac{\partial^{|\ell_{1}|}E}{\partial b^{\ell_{1}}}({\bm{x}};{\bm{A}}_{j}^{*},{\bm{b}}_{j}^{*})\frac{\partial^{|\alpha_{3}|}\mathcal{E}}{\partial{\bm{\eta}}^{\alpha_{3}}}({\bm{x}};{\bm{\eta}}_{j}^{*})$
@@
$\inf_{G\in\mathcal{M}_{N}(\Theta):\mathcal{L}_{1}(G,G_{*})\leq\varepsilon^{\prime}}\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{L}_{1}(G,G_{*})>0.$
@@
$\inf_{G\in\mathcal{M}_{N}(\Theta):\mathcal{L}_{1}(G,G_{*})>\varepsilon^{\prime}}\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{L}_{1}(G,G_{*})>0.$ (52)
Assume by contrary that the inequality (52) does not hold true, then we can find a sequence of mixing measures $G^{\prime}_{n}\in\mathcal{M}_{N}(\Theta)$ such that $\mathcal{L}_{1}(G^{\prime}_{n},G_{*})>\varepsilon^{\prime}$ and